(5.1) Find the vector form of the equation of the plane that passes through the point P0 = (1, −2, 3)

and has normal vector ~n =< 3, 1, −1 >.

(5.2) Find an equation for the plane that contains the line x = −1 + 3t, y = 5 + 3t, z = 2 + t and is

parallel to the line of intersection of the planes x −2(y −1) + 3z = −1 and y = −2x −1 = 0.


1
Expert's answer
2021-08-16T11:14:48-0400

Question 5.1 :

The points through which the plane passing is,

P0 = (1, −2, 3)

Normal vector to the plane is,

n =< 3, 1, −1 >


The vector equation of plane is given by,


"(\overrightarrow{r}-\overrightarrow{a}).\overrightarrow{n}=0"

"\overrightarrow{r}=x\overrightarrow{i}+y\overrightarrow{j}+z\overrightarrow{k}" is the position vector for any point (x,y,z) in the plane.

"\overrightarrow{a}" is the position vector of point through which plane passing "\overrightarrow{n}" is the normal vector to the plane.


Position vector of point P0(1,-2,3) is,

"\overrightarrow{a}=\hat{i}-2\hat{j}+3\hat{k}\\\overrightarrow{n}=3\hat{i}+\hat{j}-\hat{k}"


the vector equation of plane is,

"(\overrightarrow{r}-\hat{i}-2\hat{j}+3\hat{k})).(3\hat{i}+\hat{j}-\hat{k})"


this can also be written as,

"[(x\hat{i}+y\hat{j}+z\hat{k})-(\hat{i}-2\hat{j}+3\hat{k})].(3\hat{i}+\hat{j}-\hat{k})\\=[(x-1)\hat{i}+(y+2)\hat{j}+(z-3)\hat{k}].(3\hat{i}+\hat{j}-\hat{k})"



by taking the dot product of above equation, we will obtain the Cartesian form of equation of plane.


2)

"x−2(y−1)+3z=−1\\
y = −2x −1\\
y=-2x-1\\
x+4x+4+3z=-1\\
y=-2x-1\\
z=-\dfrac{5}{3}x-\dfrac{5}{3}\\
<0,−1,−\frac{5}{3}>+t<1,−2,−\frac{5}{3}>\\

The\ plane\ contains\ the\ line\\
 x = −1 + 3t, y = 5 + 3t, z = 2 + t\\

t=0: Point(-1, 5, 2)\\


t=-2: Point(-7, -1, 0)\\



-a+5b+2c+d=0\\
-7a-b+d=0\\


b=-7a+d\\

-a-35a+5d+2c+d=0\\


b=-7a+d\\
c=18a-3d\\
Substitute\\



a+14a-2d-30a+5d=0\\
d=5a\\
If\ a=1\\



d=5\\
b=-2\\
c=3\\
The\ equation\ for\ the\ plane\ is\\



x-2y+3z+5=0"